# MCDFN: Supply Chain Demand Forecasting via an Explainable Multi-Channel Data Fusion Network Model

### v2

@@ Abstract
Accurate demand forecasting is vital for optimizing supply chain management and enhancing organizational resilience. Traditional forecasting methods, relying on simple arithmetic, often fail to capture complex patterns caused by seasonal variability and special events. Although deep learning techniques have advanced, the lack of interpretable models hampers understanding and explaining predictions. We introduce the Multi-Channel Data Fusion Network (MCDFN), a novel hybrid deep learning architecture integrating multiple data modalities for superior demand forecasting. MCDFN utilizes Convolutional Neural Networks, Long Short-Term Memory networks, and Gated Recurrent Units to extract spatial and temporal features from time series data. Extensive benchmarking against seven other deep-learning models validates MCDFN’s efficacy, showing it outperforms its counterparts across key metrics with a mean squared error of 23.5738%, root mean squared error of 4.8553%, mean absolute error of 3.9991%, and mean absolute percentage error of 20.1575%. A paired t-test with a p-value of 5% indicated no significant difference between MCDFN’s predictions and actual values. A 10-fold cross-validated statistical paired t-test further confirmed its robustness, significantly outperforming other models at a 5% significance level. To address the ”black box” nature of MCDFN, we employ explainable AI techniques such as ShapTime and Permutation Feature Importance, offering insights into model decision-making processes. This research advances demand forecasting methodologies and provides practical guidelines for integrating MCDFN into existing supply chain systems while also outlining future research directions to enhance scalability, integrate external factors for better predictions, and develop user-friendly interfaces for real-world applications.
@@ 1 Introduction
Despite significant advancements in the application of DL and ML for demand forecasting in SCM, several research gaps persist that we aim to address in this paper. Specifically, there is a pressing need for more robust and scalable DL models tailored to handle the complexities and dynamic nature of SCs. Existing models often encounter challenges related to scalability and robustness, particularly when applied to large, real-world datasets. Although the current research trend in AI has witnessed significant improvements with the emergence of generative AI, the utilization of hybrid and recent interpretable DL models remains largely unexplored in the domain of SC demand forecasting. Moreover, current DL models often lack full integration of external variables such as economic indicators, weather conditions, and special events, which are known to have a significant impact on demand. Incorporating domain-specific knowledge and external factors can enhance the predictive power of these models. Another significant challenge with DL models is their inherent ”black box” nature, which can impede trust and adoption among practitioners. Therefore, enhancing the interpretability and transparency of these models is crucial for facilitating their practical utility. Although hybrid models that combine different DL techniques have shown promise, there is still ample room for improvement in their application to multi-horizon forecasting tasks. By addressing these gaps, this research endeavors to advance the state of DL in demand forecasting, ultimately providing more accurate, interpretable, and actionable insights for SCM practitioners.
@@
The paper is organized as follows: In “Literature Review” section, we reviewed some previous works related to DL and forecasting. The forecasting techniques discussed include both DL and non-DL-based methods. The “Methodology” section includes the methodology related to data collection, preprocessing, feature engineering, experimental setup, model architecture, training, and hyperparameter optimization. In the “Results” section, we showed the comparative performance of the implemented models based on the benchmarking criteria. We also validated our proposed model’s robustness using statistical paired t-tests and interpreted the model’s predictions. The “Discussions” section presents our discussions on the practical and managerial implications of our research along with the benefits and significance of our proposed model. The “Conclusions” section summarized the whole research, discussed the limitations of this research, future scopes of development, and concluded the research.
@@ 3.2.2 CNN
A dense layer with 192 units is introduced to enable non-linear transformations and higher-level feature extraction. Finally, the output layer, consisting of a dense layer with 30 units, predicts the target values for the subsequent 30 time steps. A Reshape layer is applied to reshape the output into the desired format, aligning with the temporal structure of the forecasting task. The Hyperband tuner selected the above optimal hyperparameters when the model demonstrated a minimized validation mean squared error of 26.23% after 85 trials.
@@ 3.2.3 RNN
A Dropout layer, with a dropout rate tunable from 0 to 0.5 in steps of 0.1, is included to mitigate overfitting. The model then flattens the RNN outputs and passes them through a Dense layer initialized with zeros, followed by a reshaping layer to ensure the output dimensions match the desired forecast horizon of the next 30 days. The Hyperband tuner selected the optimal recurrent units of 128 neurons and dropout of 0.1 for RNN when the model demonstrated a minimum validation mean squared error of 23.37% after 84 trials.
@@ 3.2.5 Stacked LSTM
The stacked LSTM model was designed using TensorFlow’s Keras API, which begins with an input layer for data shaped $(30,10)$, followed by two LSTM layers. Each LSTM layer is configured with 512 units, which was determined through hyperparameter tuning and employs ‘tanh’ activation and ‘sigmoid’ recurrent activation functions. Both LSTM layers use the ‘return_sequences=True’ parameter to pass the sequence output to the next layer. Between the LSTM layers, a Dropout layer is included to mitigate overfitting, with 0 dropout rate after tuning. After the LSTM layers, the model output is flattened and passed through a Dense layer, initialized with zeros, ensuring unbiased predictions. Finally, the output is reshaped to the desired dimensions (30, 1). Stacked LSTM was evaluated and achieved the best validation mean squared error of 28.17% after 58 trials, demonstrating its ability to capture complex temporal dependencies and provide accurate demand forecasts for an output sequence length of 30 steps.
@@ 3.2.6 MLP
Our research utilized a Multilayer Perceptron (MLP) model for demand forecasting, implemented using TensorFlow’s Keras API. The model starts with an input layer to handle data shaped $(30,10)$, followed by a dense layer with 512 units chosen through hyperparameter tuning. The activation function for this layer was selected as ‘tanh’ from the options ‘relu’ and ‘tanh’. A Dropout layer with a rate of 0.0 is included to prevent overfitting. Finally, the model outputs predictions through a dense layer with a linear activation function to match the dimensionality of the target feature. The MLP model was evaluated and achieved the best validation mean squared error of 27.79% after 54 trials.
@@ 3.2.7 GRU
Our research employed a Gated Recurrent Unit (GRU) model, which begins with an input layer to handle data shaped $(30,10)$, followed by a GRU layer configured with 192 units, a hyperparameter value determined through extensive tuning. The GRU layer uses ‘return_sequences=True’ to maintain the sequential nature of the data. A Dropout layer with a dropout rate of 0.4 is included to prevent overfitting. After the GRU and dropout layers, the model’s output is flattened and passed through a Dense layer, initialized with zeros, to ensure unbiased predictions. Finally, the output is reshaped to match the required dimensions of (30, 1). GRU model achieved the best validation mean squared error of 23.84% during tuning after 16 trials, demonstrating its capability to accurately capture temporal dependencies and provide robust demand forecasts for an output sequence length of 30 steps. This architecture and its hyperparameters were optimized to enhance prediction accuracy while managing model complexity.
@@ 3.2.8 Proposed MCDFN Model
We used hyperparameter tuning to optimize the number of filters in the CNN layers, kernel sizes, and units in the LSTM and GRU layers. After 90 trials of hyperparameter search, the optimal CNN filters, kernel size, and LSTM/GRU units were found to be 352, 1, and 64, respectively, in the 16th trial. The model with these hyperparameters achieved the lowest validation MSE of 23.29% during trials and the tuning process.
@@ 4.2 Comparative Analysis of Results
Model Sample of data Loss (%) MSE (%) RMSE (%) MAE (%) MAPE (%)
@@ 6 Conclusions
We validated the robustness and reliability of MCDFN through a 10-fold cross-validated statistical paired t-test for each benchmarked model’s evaluation metrics, showing that MCDFN significantly outperforms other models at a 5% significance level. Additionally, we performed a paired t-test for each model using a p-value of 5%, demonstrating no significant difference between the actual and predicted values of MCDFN. To address the “black box” nature of MCDFN, we employed explainable artificial intelligence (XAI) techniques such as ShapTime and PFI, providing insights into the model’s decision-making process.
@@
CRediT Author Statement
Md Abrar Jahin: Writing – Original draft preparation, Conceptualization, Methodology, Software, Formal analysis, and Visualization. Asef Shahriar: Writing – Original draft preparation, Conceptualization, Methodology, Software, and Data curation. Md Al Amin: Supervision, Writing – Review & Editing.

### v3

@@ Abstract
Accurate demand forecasting is vital for optimizing supply chain management and enhancing organizational resilience. Traditional forecasting methods, relying on simple arithmetic, often fail to capture complex patterns caused by seasonal variability and special events. Although deep learning techniques have advanced, the lack of interpretable models hampers understanding and explaining predictions. We introduce the Multi-Channel Data Fusion Network (MCDFN), a novel hybrid deep learning architecture integrating multiple data modalities for superior demand forecasting. MCDFN utilizes Convolutional Neural Networks, Long Short-Term Memory networks, and Gated Recurrent Units to extract spatial and temporal features from time series data. Comparative benchmarking against seven other deep-learning models validates MCDFN’s efficacy, showing it outperforms its counterparts across key metrics with a mean squared error of 23.5738, root mean squared error of 4.8553, mean absolute error of 3.9991, and mean absolute percentage error of 20.1575%. A paired t-test with a p-value of 5% indicated no significant difference between MCDFN’s predictions and actual values. A 10-fold cross-validated statistical paired t-test further confirmed its robustness, significantly outperforming other models at a 5% significance level. To address the “black box” nature of MCDFN, we employ explainable AI techniques such as ShapTime and Permutation Feature Importance, offering insights into model decision-making processes. This research advances demand forecasting methodologies and provides practical guidelines for integrating MCDFN into existing supply chain systems while also outlining future research directions to enhance scalability, integrate external factors for better predictions, and develop user-friendly interfaces for real-world applications.
@@ 1 Introduction
Despite significant advancements in the application of DL and ML for demand forecasting in SCM, several research gaps persist that we aim to address in this paper. Specifically, there is a pressing need for more robust and scalable DL models tailored to handle the complexities and dynamic nature of SCs. Existing models often encounter challenges related to scalability and robustness, particularly when applied to large, real-world datasets. Although the current research trend in AI has witnessed significant improvements with the emergence of generative AI, the utilization of hybrid and recent interpretable DL models remains largely unexplored in the domain of SC demand forecasting. Moreover, current DL models often lack full integration of external variables such as economic indicators, weather conditions, and special events, which are known to have a significant impact on demand. Incorporating domain-specific knowledge and external factors can enhance the predictive power of these models. Another significant challenge with DL models is their inherent “black box” nature, which can impede trust and adoption among practitioners. Therefore, enhancing the interpretability and transparency of these models is crucial for facilitating their practical utility. Although hybrid models that combine different DL techniques have shown promise, there is still ample room for improvement in their application to multi-horizon forecasting tasks. By addressing these gaps, this research endeavors to advance the state of DL in demand forecasting, ultimately providing more accurate, interpretable, and actionable insights for SCM practitioners.
@@
The paper is organized as follows: In “Literature Review” section, we reviewed some previous works related to DL and forecasting. The forecasting techniques discussed include both DL and non-DL-based methods. The “Methodology” section includes the methodology related to data collection, preprocessing, feature engineering, experimental setup, model architecture, training, and hyperparameter optimization. In the “Results” section, we showed the comparative performance of the implemented models based on the benchmarking criteria. We also validated our proposed model’s robustness using statistical paired t-tests and interpreted the model’s predictions. The “Discussion” section presents our discussions on the practical and managerial implications of our research along with the benefits and significance of our proposed model. The “Conclusions” section summarized the whole research, discussed the limitations of this research, future scopes of development, and concluded the research.
@@ 3.2.2 CNN
A dense layer with 192 units is introduced to enable non-linear transformations and higher-level feature extraction. Finally, the output layer, consisting of a dense layer with 30 units, predicts the target values for the subsequent 30 time steps. A Reshape layer is applied to reshape the output into the desired format, aligning with the temporal structure of the forecasting task. The Hyperband tuner selected the above optimal hyperparameters when the model demonstrated a minimized validation mean squared error of 26.23 after 85 trials.
@@ 3.2.3 RNN
A Dropout layer, with a dropout rate tunable from 0 to 0.5 in steps of 0.1, is included to mitigate overfitting. The model then flattens the RNN outputs and passes them through a Dense layer initialized with zeros, followed by a reshaping layer to ensure the output dimensions match the desired forecast horizon of the next 30 days. The Hyperband tuner selected the optimal recurrent units of 128 neurons and dropout of 0.1 for RNN when the model demonstrated a minimum validation mean squared error of 23.37 after 84 trials.
@@ 3.2.5 Stacked LSTM
The stacked LSTM model was designed using TensorFlow’s Keras API, which begins with an input layer for data shaped $(30,10)$, followed by two LSTM layers. Each LSTM layer is configured with 512 units, which was determined through hyperparameter tuning and employs ‘tanh’ activation and ‘sigmoid’ recurrent activation functions. Both LSTM layers use the ‘return_sequences=True’ parameter to pass the sequence output to the next layer. Between the LSTM layers, a Dropout layer is included to mitigate overfitting, with 0 dropout rate after tuning. After the LSTM layers, the model output is flattened and passed through a Dense layer, initialized with zeros, ensuring unbiased predictions. Finally, the output is reshaped to the desired dimensions (30, 1). Stacked LSTM was evaluated and achieved the best validation mean squared error of 28.17 after 58 trials, demonstrating its ability to capture complex temporal dependencies and provide accurate demand forecasts for an output sequence length of 30 steps.
@@ 3.2.6 MLP
Our research utilized a Multilayer Perceptron (MLP) model for demand forecasting, implemented using TensorFlow’s Keras API. The model starts with an input layer to handle data shaped $(30,10)$, followed by a dense layer with 512 units chosen through hyperparameter tuning. The activation function for this layer was selected as ‘tanh’ from the options ‘relu’ and ‘tanh’. A Dropout layer with a rate of 0.0 is included to prevent overfitting. Finally, the model outputs predictions through a dense layer with a linear activation function to match the dimensionality of the target feature. The MLP model was evaluated and achieved the best validation mean squared error of 27.79 after 54 trials.
@@ 3.2.7 GRU
Our research employed a Gated Recurrent Unit (GRU) model, which begins with an input layer to handle data shaped $(30,10)$, followed by a GRU layer configured with 192 units, a hyperparameter value determined through extensive tuning. The GRU layer uses ‘return_sequences=True’ to maintain the sequential nature of the data. A Dropout layer with a dropout rate of 0.4 is included to prevent overfitting. After the GRU and dropout layers, the model’s output is flattened and passed through a Dense layer, initialized with zeros, to ensure unbiased predictions. Finally, the output is reshaped to match the required dimensions of (30, 1). GRU model achieved the best validation mean squared error of 23.84 during tuning after 16 trials, demonstrating its capability to accurately capture temporal dependencies and provide robust demand forecasts for an output sequence length of 30 steps. This architecture and its hyperparameters were optimized to enhance prediction accuracy while managing model complexity.
@@ 3.2.8 Proposed MCDFN Model
We used hyperparameter tuning to optimize the number of filters in the CNN layers, kernel sizes, and units in the LSTM and GRU layers. After 90 trials of hyperparameter search, the optimal CNN filters, kernel size, and LSTM/GRU units were found to be 352, 1, and 64, respectively, in the 16th trial. The model with these hyperparameters achieved the lowest validation MSE of 23.29 during trials and the tuning process.
@@ 4.2 Comparative Analysis of Results
Model Sample of data Loss MSE RMSE MAE MAPE (%)
@@ 6 Conclusions
We validated the robustness and reliability of MCDFN through a 10-fold cross-validated statistical paired t-test for each benchmarked model’s evaluation metrics, showing that MCDFN significantly outperforms other models at a 5% significance level. Additionally, we performed a paired t-test for each model using a p-value of 5%, demonstrating no significant difference between the actual and predicted values of MCDFN. To address the “black box” nature of MCDFN, we employed XAI techniques such as ShapTime and PFI, providing insights into the model’s decision-making process.